# arXiv:2603.01211v1[cs.AI] 01 Mar 2026

### v1

@@ Abstract
As generative AI technologies are increasingly being launched across the globe, assessing their competence to operate in different cultural contexts is exigently becoming a priority. While recent years have seen numerous and much-needed efforts on cultural benchmarking, these efforts have largely focused on specific aspects of culture and evaluation. While these efforts contribute to our understanding of cultural competence, a unified and systematic evaluation approach is needed for us as a field to comprehensively assess diverse cultural dimensions at scale. Drawing on measurement theory, we present a principled framework to aggregate multifaceted indicators of cultural capabilities into a unified assessment of cultural intelligence. We start by developing a working definition of culture that includes identifying core domains of culture. We then introduce a broad-purpose, systematic, and extensible framework for assessing cultural intelligence of AI systems. Drawing on theoretical framing from psychometric measurement validity theory, we decouple the background concept (i.e., cultural intelligence) from its operationalization via measurement. We conceptualize cultural intelligence as a suite of core capabilities spanning diverse domains, which we then operationalize through a set of indicators designed for reliable measurement. Finally, we identify the considerations, challenges, and research pathways to meaningfully measure these indicators, specifically focusing on data collection, probing strategies, and evaluation metrics.
@@ 1 Introduction
Artificial intelligence (AI) has rapidly transitioned from a niche technology to a ubiquitous component of our daily lives. Millions now interact with generative AI systems for applications ranging from seeking information [72], to complex problem-solving [26], to companionship [52], while specialized AI models are reshaping core knowledge ecosystems in domains such as health [98], education [103], and art [100]. These advancements are also accelerating from generative tools toward more autonomous agentic tools [3, 73], with capabilities claimed to be approaching general intelligence [10]. While the benchmark-driven development [13, 82] has been instrumental in advancing AI performance in logic, math, and language, existing evaluations often overlook whether a model can truly navigate the contextual sensitivities inherent to diverse global cultures. This gap becomes critical as these models are increasingly deployed in social, and often sensitive, roles such as companionship and education. Models can exhibit a superficial, yet convincing, illusion of cultural intelligence, for instance, by conversing fluently in a local language, or responding to certain culture-laden requests, but without specific and relevant cultural knowledge, thus creating a false sense of competency and safety for users across cultural contexts.
@@
By “cultural intelligence”, we refer to a model’s capacity to strategically adapt across diverse contexts, generating responses that are not only factually grounded in accurate cultural knowledge but also respectful of relevant norms and values (we define this term formally in Section 2.2). Failing to assess and improve cultural intelligence in AI may give rise to a multitude of risks. The first is a functional risk: models cannot be truly helpful unless they are competent to serve user needs around local cultural knowledge in local languages they may prefer, while also adhering to implicit local norms (e.g., [42, 101]). This may lead directly to both a socio-economic risk—as performance disparities across cultural contexts will limit AI proliferation, adoption, and economic benefit in new markets [16], as well as an equity risk—erecting barriers for diverse communities from accessing the benefits of technological advancements [8]. There is also the critical safety risk: culturally incompetent models may overlook violations of social norms and cause unique safety failures preventable through a cultural lens [91]. Finally, and perhaps most profoundly, there is the systemic ethical risk of reproducing a dominant, hegemonic culture, that leads to the exclusion and marginalization of other worldviews, leading to value imposition and cultural erasure [62, 54].
@@
Some of these risks and failures have already been empirically demonstrated, ranging from the inability to accurately answer culturally specific general knowledge questions [55, 81] or depict salient artifacts [43], to critical failures in adhering to safety guardrails [83, 30], while reinforcing dominant sub-cultures [67] and propagating harmful cultural stereotypes [53]. There have also been a surge in benchmarks addressing these issues, for instance, measuring “everyday knowledge” (e.g., BLEnD [55], Global MMLU [81]), and situational etiquette (e.g., NORMAD [69]). Empirical efforts have also probed values encoded in LLMs [2] and vision-language models [94], while [36] assess how LLMs adapt their linguistic style to match cultural expectations. While these works are vital, they often reduce culture to isolated data points [59], and tend to address the problem in a piecemeal fashion, examining individual components of the puzzle rather than treating cultural intelligence as a cohesive whole. While recent surveys (e.g., [48, 88]) offer useful categorizations of these efforts, they often fail to provide a unifying and systematic evaluation framework for conceptualizing and synthesizing the diverse capabilities that contribute to cultural intelligence in AI, nor do they engage with the inherent challenges in that endeavor. Such a framework is imperative not only for tracking progress on integrating these capabilities into the pursuit of AGI, but also for broader stakeholders such as policy makers and governance bodies to comprehensively assess AI capabilities and safety across diverse cultural contexts.
@@
Addressing this need raises a core question: how should cultural capabilities of AI systems be conceptualized, measured, and improved in ways which illuminate the progress of the AI field towards intelligence more generally? There is a multitude of reasons why this remains a grand, open, research challenge in the AI community. First, there currently is a lack of an operational understanding or definition of culture in AI literature; “culture” is a complex and contested term to define and quantify [89]. Furthermore, culture is inherently pluralistic: societies differ profoundly not only in their artifacts, practices, and norms, but also in their interpretations, meanings, and associations. They also differ in salience, as the relative importance of specific aspects like religion or rituals varies dramatically across cultures. Secondly, even with a working definition of culture, we currently lack a systematic approach towards quantifying and evaluating cultural intelligence. A comprehensive evaluation of cultural intelligence should consider its various facets, across diverse use cases that end-users use AI models for. Finally, even with a comprehensive set of evaluation criteria, we lack broad-coverage and high-quality data that can ground and enable large-scale evaluation and improvement of cultural intelligence, across locales, languages, and modalities.
@@
In this paper, we introduce a broad-purpose, systematic framework for assessing cultural intelligence of AI systems. Adopting the four-level measurement approach of Adcock and Collier [1] rooted in psychometric measurement validity theory, we explicitly decouple the background concept (i.e., cultural intelligence) from its operationalization via measurement. We structure this process in three steps. Conceptualization: to establish a clear basis for measurement, we first develop a systematized concept of cultural intelligence, defined in terms of a set of core capabilities that span different cultural domains. Operationalization: we then operationalize the concept of cultural intelligence as a set of indicators designed to reliably measure these capabilities collectively. Measurement: finally, we identify the considerations, challenges, and research pathways to meaningfully measure these indicators, specifically focusing on data collection, probing strategies, and evaluation metrics. This structured and fine-grained framework transforms cultural intelligence into a measurable construct that is extensible and controllable in its application across diverse contexts, while also interpretable to generate actionable insights.
@@ 2.1 Defining Culture for AI: A Pragmatic Ontology
Culture is a complex and contested term, with definitions that have shifted significantly over time across anthropology, sociology, and history [89]. Rather than navigating these fluctuating definitions or attempting to define culture in the abstract, we adopt a pragmatic approach [77], conceptualizing culture toward the operational goal of assessing and improving the cultural intelligence of AI systems.
As a foundation, we develop a Vocabulary of Culture organized around three high-level domains that appear consistently across theoretical work: Cultural Production, Behavior and Practices, and Knowledge and Values. This triad reflects a recurring idea in social science attempts to define culture: Bodley [14, p22] summarizes culture as “what people think, what they do, and the material products they produce,” while Spradley [80] similarly distinguishes between cultural artifacts, cultural behavior, and the implicit cultural knowledge used to interpret experience. While some traditional definitions often split culture simply into material (tangible) and non-material (intangible) aspects, AI assessment benefits from the finer tripartite distinction. In particular, we must distinguish human internal determinants, such as beliefs, values, and norms, from behavioral outputs like rituals and practices; evaluating a model’s latent knowledge differs fundamentally from evaluating its generated actions. Consequently, our Vocabulary of Culture (Table 1) categorizes cultural aspects into:
@@
  • Knowledge and Values or how people think, believe or know, such as moral codes, shared meanings, and collective associations such as stereotypes.
@@
Table 1 identifies some finer-grained topics within each domain and some specific facets within each topic. It further lists an exemplar set of relevant literature which has contributed theory, data or evaluations around the specific topic. While this ontology was constructed primarily using thought experiments over geo-cultures, we posit that it is adaptable to cultures defined by religion, language, or other sociological groupings. It is important to note that this list is illustrative rather than exhaustive. Certain facets may bridge categories; for instance, sports can be analyzed as both a practiced event and an outcome of cultural production.
@@ 2.1.1 Methodology of Vocabulary Construction
To develop a robust vocabulary of culture, we employed a dual-stream methodology that synthesized bottom-up empirical observations with top-down theoretical frameworks. Going bottom-up, we first curated an initial list of cultural elements by auditing prominent benchmarks within the Machine Learning, Natural Language Processing, and Computer Vision literature. This included both domain-specific cultural datasets [57, 55, 93] and general-purpose benchmarks. The objective was to identify downstream, human-centered tasks where cultural nuances, whether explicit or latent could potentially impact model performance. These disparate elements were then clustered into preliminary categories and iteratively expanded using LLM-assisted augmentation to identify potential sub-categories. Simultaneously going top-down, we also conducted a literature review across anthropology, sociology, and linguistics to understand culture as seen by society and associated studies. This approach ensured that the vocabulary remained grounded in established social theory, providing a deductive framework to categorize the multifaceted nature of “culture” beyond mere data labels. This dual-stream approach enabled the performance of a horizontal reconciliation of the two respective lists generated by the approaches to resolve ontological discrepancies and formalize a comprehensive functional vocabulary. To ensure face validity and interdisciplinary coverage, this draft was then subjected to a rigorous feedback loop with the full authorship team, representing a diverse range of expertise. The iterative refinement resulted in the final hierarchical “Vocabulary of Culture”.
@@ 2.2 Cultural Intelligence as a set of Capabilities
If operationalizing culture is difficult, operationalizing the evaluation of cultural intelligence is an even harder task. The relationship of culture to intelligence, in particular, has been complicated by the AI field’s lack of consensus around the construct of intelligence itself. To address this, we must first define what we mean by “intelligence” and “culture” in the context of AI systems. As we are primarily interested in evaluation, we adopt a definition of intelligence by Bryson [15] that anchors on a model’s external behavior (rather than internal reasoning): the capacity to act effectively in response to changing contexts. Under this definition, (artificial) intelligence should be ascribed to computer systems when they act effectively for the diverse contexts of their deployments. When these deployment contexts involve interacting with human users or communities, cultural dimensions of the context are inherited from the cultural capabilities and contexts of those users or communities. We thus propose the following working definition for cultural intelligence of AI systems:
@@ 2.2.1 Cultural Sensing
A culturally intelligent model must first possess the Cultural Sensing capability, the discriminative ability to detect when a request invokes specific cultural domains or topics in ways that should inform the response. Without this capability, a model risks treating culturally laden requests with a “neutral” or universalist lens, leading to functional failures or safety violations — for instance, recommending culturally taboo ingredients for a meal plan request because it failed to recognize the cultural weight of such a request. Cultural sensing essentially acts as a gatekeeper, ensuring the system transitions from generic processing for agnostic inputs, to a culturally situated processing for laden inputs:
@@
Cultural sensing is inherently tied to understanding the user intent, as a model cannot accurately assess “cultural weight” without understanding the specific user needs and underlying intentions. Identifying whether a user seeks factual retrieval, creative inspiration, or social navigation, will help the model determine if a cultural lens is required. Taxonomies of usage, such as those proposed in recent literature (e.g., [74]), provide instrumental “intent buckets” that could help parse the cultural weight of the request. For example, an intent of seeking social advice may proactively trigger sensing for cultural norms, whereas a mathematical calculation might safely bypass the cultural lens as agnostic.
@@ 2.2.2 Cultural Scoping
Once a cultural signal is sensed, the model should have the capability to discern the appropriate cultural scope (e.g., Global vs. Locale Specific) that a user query is about and the specific cultural context that the response should acknowledge and invoke. It needs to be able to map the query to the specific cultural group relying on:
@@
  • implicit cues — for instance, mention of cultural artifacts (e.g., mention of “Poriyal” may help resolve the scope to India or Tamil Nadu, more specifically), dialectal terms (e.g., “flat” to refer to a residence), or the language used in the request.
@@
With these In the absence of specific cues, the model should acknowledge the ambiguity appropriately in its response, and avoid Western defaults or other biases. This phase inherently manages noisy, ambiguous, or contradictory signals—for instance, the dissonance between a US-based IP address and the dialectal use of “flat.” Consequently, the system requires a strategy to calibrate the confidence of its scoping assessment. Even when a definitive scope cannot be established, partial information must be preserved as a critical input for downstream processing.
@@ 2.2.3 Cultural Fluency
Once the cultural scope is established, the core capability is Cultural Fluency — the ability to construct a culturally fluent output. We identify three core components to assess cultural fluency:
  • Epistemic Fidelity (The Knowledge): The capability of the model to retrieve cultural knowledge that are both comprehensive (high coverage across cultural artifacts) and accurate (being able to answer factual questions about cultural artifacts correctly). This includes identifying the correct ingredients in a food item, or the correct steps in a ritual, or the correct norms in a scenario, given a cultural context.
  • Representational Richness (The Diversity): The capability to determine and represent richness within the cultural scope, representing diversity and conceptual breadth, avoiding simplification. This could include the models presenting a response with global diversity for a globally scoped query (e.g., tortilla maps to different items in Spain vs. Mexico) or non-material artifacts (e.g., norms around tipping varies across cultures).
  • Pragmatic Proficiency (The Application): The skillful application of cultural knowledge, norms, and values to ensure the response is socially congruent. This may include both form and content — applying the cultural knowledge within the conversation (e.g., following relevant linguistic norms) and reasoning with the relevant cultural knowledge in generating responses (e.g., problem solving for the user in ways that don’t violate cultural norms).
The cultural fluency capabilities are in an ascending order of complexity, somewhat similar to Bloom’s taxonomy [11]. While epistemic fidelity primarily requires ‘remembering’ of information within the model, representational richness requires a deeper ‘understanding’ and compilation of knowledge, and finally pragmatic proficiency requires the ability to use, analyze, and even create new information. Each capability thus can consist of different types of actions and can be evaluated for by leveraging different tasks that demonstrate said capability either partially or completely. Table 3 provides a set of examples that trace how our framework will assess different scenarios, and what the desirable behavior of a culturally intelligent model should be. While any single query may evoke the need for multiple cultural intelligence capabilities, for simplicity, we choose to highlight only one of the capabilities for each example.

### v2

@@ Abstract
As generative AI technologies are increasingly being launched across the globe, assessing their competence to operate in different cultural contexts is urgently becoming a priority. While recent benchmarking efforts have made revealed significant gaps in cultural competence of AI models, they remain fragmented, often focusing on specific cultural aspects or evaluation tasks. This fragmentation hinders the ability of model developers, users, and policymakers to comprehensively assess the readiness of these models to be deployed or used in specific regional or cultural environments. To bridge this gap, we propose a unified, structured, and fine-grained framework for assessing cultural intelligence at scale. Drawing on psychometric measurement theory, our approach decouples the conceptual definition of cultural intelligence from its operationalization. We start by developing a working definition of culture that includes identifying core domains, topics, and facets of culture. We then conceptualize cultural intelligence as a suite of latent capabilities spanning diverse domains, and demonstrate operationalizing these capabilities through a series of measurable indicators designed for reliability and extensibility. Finally, we identify the practical considerations and research pathways for data and metrics, providing a scalable roadmap for the comprehensive and responsible assessment of AI cultural competence.
@@ 1 Introduction
Artificial intelligence (AI) has rapidly transitioned from a niche technology to a ubiquitous component of our daily lives. Millions now interact with generative AI systems for applications ranging from seeking information [1], to complex problem-solving [2], to companionship [3], while specialized AI models are reshaping core knowledge ecosystems in domains such as health [4], education [5], and art [6]. These advancements are also accelerating from generative tools toward more autonomous agentic tools [7, 8], with capabilities claimed to be approaching general intelligence [9]. While the benchmark-driven development [10, 11] has been instrumental in advancing AI performance, especially in logic, math, language, and reasoning, existing evaluations often overlook whether a model can truly navigate the contextual sensitivities inherent to diverse global cultures. This gap becomes critical as these models are increasingly deployed in social, and often sensitive, roles such as companionship and education. Models can exhibit a superficial, yet convincing, illusion of cultural intelligence, for instance, by conversing fluently in a local language, or responding to certain culture-laden requests, but without reliable culturally grounded reasoning, thus creating a false sense of competency and safety for users across cultural contexts.
@@
By “cultural intelligence”, we refer to a model’s capacity to strategically adapt across diverse contexts, generating responses that are not only factually grounded in accurate cultural knowledge but also respectful of relevant norms and values (we define this term formally in Section 2.2). Failing to assess and improve cultural intelligence in AI may give rise to a multitude of risks. The first is a functional risk: models cannot be truly helpful unless they are competent to serve user needs around local cultural knowledge in local languages they may prefer, while also adhering to implicit local norms (e.g., [12, 13]). This may lead directly to both a socio-economic risk—as performance disparities across cultural contexts will limit AI proliferation, adoption, and economic benefit in new markets [14], as well as an equity risk—erecting barriers for diverse communities from accessing the benefits of technological advancements [15]. There is also the critical safety risk: culturally incompetent models may cause unique safety failures where standard guardrails break down, overlooking offensive violations of social norms that are only detectable through a nuanced cultural lens [16]. Finally, and perhaps most profoundly, there is the systemic ethical risk of reproducing a dominant, hegemonic culture, that leads to the exclusion and marginalization of other worldviews, leading to value imposition and cultural erasure [17, 18].
@@
Some of these risks and failures have already been empirically demonstrated, ranging from the inability to accurately answer culturally specific general knowledge questions [19, 20] or depict salient artifacts [21], to critical failures in adhering to safety guardrails [22, 23], while reinforcing dominant sub-cultures [24] and propagating harmful cultural stereotypes [25]. Recent benchmarks attempt to address these issues, for instance, measuring linguistic style [26], “everyday knowledge” (e.g., BLEnD [19], Global MMLU [20]), and situational etiquette (e.g., NORMAD [27]). Empirical efforts have also probed values encoded in LLMs [28] and vision-language models [29]. While these works are vital, they often reduce culture to isolated data points [30], and tend to address the problem in a piecemeal fashion, examining individual components of the puzzle rather than treating cultural intelligence as a cohesive whole. While recent surveys (e.g., [31, 32]) offer useful categorizations of these efforts, they often fail to provide a unifying and systematic evaluation framework for conceptualizing and synthesizing the diverse capabilities that contribute to cultural intelligence in AI, nor do they engage with the inherent challenges in that endeavor. Such a framework is imperative not only for tracking progress on integrating these capabilities into the pursuit of AGI, but also for broader stakeholders such as policy makers and governance bodies to comprehensively assess AI capabilities and safety across within local cultural contexts.
@@
Addressing this need raises a core question: how should cultural capabilities of AI systems be conceptualized, measured, and improved? There is a multitude of reasons why this remains a grand, open, research challenge in the AI community. First, there currently is a lack of an operational understanding or definition of culture in AI literature; “culture” is a complex and contested term to define and quantify [33]. Furthermore, culture is inherently pluralistic: societies differ profoundly not only in their artifacts, practices, and norms, but also in their interpretations, meanings, and associations. They also differ in salience, as the relative importance of specific aspects like religion or rituals varies dramatically across cultures. Secondly, even with a working definition of culture, we currently lack a systematic approach towards quantifying and evaluating cultural intelligence. A comprehensive evaluation of cultural intelligence should consider its various facets, across diverse use cases that end-users use AI models for. Finally, even with a comprehensive set of evaluation criteria, we lack broad-coverage and high-quality data that can ground and enable large-scale evaluation and improvement of cultural intelligence, across locales, languages, and modalities.
@@
In this paper, we introduce a broad-purpose, structured, and fine-grained framework for assessing cultural intelligence of AI systems. Adopting the four-level measurement approach of Adcock and Collier [34] rooted in psychometric measurement validity theory, we explicitly decouple the background concept (i.e., cultural intelligence) from its operationalization. We structure this process in three steps. Conceptualization: to establish a clear basis for measurement, we first develop a systematized concept of cultural intelligence, defined in terms of a set of core capabilities that span different cultural domains. Operationalization: we then operationalize each of these capabilities as an extensible set of indicators designed to reliably measure cultural intelligence, collectively. Measurement: finally, we identify the considerations, challenges, and research pathways to meaningfully measure these indicators, specifically focusing on data collection, probing strategies, and evaluation metrics. This structured and fine-grained framework transforms cultural intelligence into a measurable construct that is extensible and controllable in its application across diverse contexts, while also interpretable to generate actionable insights.
@@ 2.1 Defining Culture for AI: A Pragmatic Ontology
Culture is a complex and contested term, with definitions that have shifted significantly over time across anthropology, sociology, and history [33]. We sidestep this abstract definitional volatility in favor of a pragmatic approach [36], conceptualizing culture through a functional lens, focusing on the operational requirements for evaluating and improving the cultural intelligence of AI systems. As a foundation, we develop a hierarchical and multi-dimensional Vocabulary of Culture organized around three high-level domains that appear consistently across theoretical work: Cultural Production, Behavior and Practices, and Knowledge and Values. This triad reflects a recurring idea in social science attempts to define culture: Bodley [37, p22] summarizes culture as “what people think, what they do, and the material products they produce,” while Spradley [38] similarly distinguishes between cultural artifacts, cultural behavior, and the implicit cultural knowledge used to interpret experience. While some traditional definitions often split culture simply into material (tangible) and non-material (intangible) aspects, AI assessment benefits from the finer tripartite distinction. In particular, we must distinguish human internal determinants, such as beliefs, values, and norms, from behavioral outputs like rituals and practices; evaluating a model’s latent knowledge differs fundamentally from evaluating its generated actions. Consequently, our Vocabulary of Culture (Table 1) categorizes cultural aspects into:
@@
  • Knowledge and Values or how people think, believe or know, such as moral codes, shared meanings, and collective associations, including stereotypes.
@@
Table 1 identifies some finer-grained topics within each domain and some specific facets within each topic; it is important to note that this list is illustrative rather than exhaustive. The table further shows an exemplar set of relevant literature which has contributed theory, data or evaluations around the specific topic. While this ontology was constructed primarily using thought experiments over geo-cultures, we posit that it is adaptable to cultures defined by religion, language, or other sociological groupings.
@@ 2.1.1 Methodology of Vocabulary Construction
To develop this robust vocabulary of culture into various topics and facets, we employed a dual-stream methodology that synthesized bottom-up empirical observations with top-down theoretical frameworks. Going bottom-up, we first curated an initial list of cultural elements by auditing prominent benchmarks within Machine Learning literature. The objective was to identify downstream, human-centered tasks where cultural nuances, whether explicit or latent could potentially impact model performance. These disparate elements were then clustered into preliminary categories and iteratively expanded using LLM-assisted augmentation to identify potential sub-categories. Simultaneously going top-down, we conducted a literature review across anthropology, sociology, and linguistics to understand culture as seen by diverse disciplinary traditions. This approach ensured that the vocabulary remained grounded in established social theory, providing a deductive framework to categorize the multifaceted nature of “culture” beyond mere data labels. This dual-stream approach enabled the performance of a horizontal reconciliation of the two respective lists generated by the approaches to resolve ontological discrepancies and formalize a functional vocabulary. To ensure face validity and interdisciplinary coverage, this draft was then subjected to a rigorous feedback loop with a subset of the authors, representing diverse disciplinary expertise. The iterative refinement resulted in the final hierarchical “Vocabulary of Culture”.
@@ 2.2 Cultural Intelligence as a set of Capabilities
If operationalizing culture is difficult, operationalizing the evaluation of cultural intelligence is an even harder task. The relationship of culture to intelligence, in particular, has been complicated by the AI field’s lack of consensus around the construct of intelligence itself. As we are primarily interested in evaluation, we adopt a definition of intelligence by Bryson [62] that anchors on a model’s external behavior (rather than internal reasoning): the capacity to act effectively in response to changing contexts. Under this definition, (artificial) intelligence should be ascribed to computer systems when they act effectively for the diverse contexts of their deployments. When these deployment contexts involve interacting with human users or communities, cultural dimensions of the context are inherited from the cultural capabilities and contexts of those users or communities. We thus propose the following working definition for cultural intelligence of AI systems:
@@ 2.2.1 Cultural Sensing
A culturally intelligent model must first possess the Cultural Sensing capability, the discriminative ability to detect when a request invokes specific cultural domains or topics in ways that should inform the response. Without this capability, a model risks treating culturally laden requests with a “neutral” or universalist lens, leading to functional failures or safety violations—e.g., recommending culturally taboo ingredients for a meal plan request. Cultural sensing essentially acts as a gatekeeper, detecting user intent [63] and also ensuring the system transitions from generic processing for agnostic inputs, to a culturally situated processing for laden inputs:
@@ 2.2.2 Cultural Scoping
Once a cultural signal is sensed, the model should have the capability to discern the appropriate cultural scope (e.g., Global vs. Locale Specific) that a user query is about and the specific cultural context that the response should acknowledge and invoke. This step could rely on:
@@
  • implicit cues, such as language of the request, cultural artifact mentions (e.g., “Poriyal” indicating a Tamil context), or dialectal terms (e.g., “flat” for residence)
@@
In the absence of specific cues, the model should acknowledge the ambiguity appropriately in its response, and avoid Western defaults or other biases. This phase should manages noisy, ambiguous, or contradictory signals—for instance, the dissonance between a US-based IP address and the dialectal use of “flat.” Consequently, the system requires a strategy to calibrate the confidence of its scoping assessment.
@@ 2.2.3 Cultural Fluency
Once the cultural scope is established, the core capability is Cultural Fluency—the ability to construct a culturally fluent output. We identify three core components to assess cultural fluency, Epistemic Fidelity, Representational Richness, and Pragmatic Proficiency, which we define in Table 2. We note how the cultural fluency capabilities are in an ascending order of complexity, somewhat similar to Bloom’s taxonomy [64]. While epistemic fidelity primarily requires ‘remembering’ of information within the model, representational richness requires a deeper ‘understanding’ and compilation of knowledge, and finally pragmatic proficiency requires the ability to use, analyze, and even create new information. Each capability thus can consist of different types of actions and can be evaluated for by leveraging different indicators that demonstrate said capability, either partially or completely. Table 3 provides a set of examples that trace how our framework will assess different scenarios, and what the desirable behavior of a culturally intelligent model should be. While any single query may evoke the need for multiple cultural intelligence capabilities, for simplicity, we choose to highlight only one of the capabilities for each example.